Complete the missing blanks in the following table. Remember concentration is amount of material per until...

  1. Complete the missing blanks in the following table. Remember concentration is amount of material per until volume.   Concentrations can be expressed as g/volume, mols/volume, or even as percentages (amount/100ml).   The other important point is to keep your units correct
  2. Concentration

    Volume used

    Amount of material

    15mM

    45 ul

      

    200uM

    1 ml

      

    250mM

    15 ul

      

      

    125 ul

    200 nMols

      

    9 ml

    25 mMols

      

    10 ul

    5 uMols

    1M

      

    55 uMols

    50uM

      

    100 pMols

    80mM

      

    165 uMols
Answer #1

Part 1: The amounts of substance in each solution are calculated:

amount = M * V = 15 mmol / L * 10 uL * (1 L / 1x10 ^ -6 uL) = 1.5x10 ^ 8 mmol

2: 0.2 umol

3: 3.75x10 ^ 9 mmol

Part 2: The concentrations are calculated:

Concentrations = n / V = ​​(200 nmol / 125 uL) * (1x10 ^ -6 uL / ​​1 L) = 1.6x10 ^ 6 nM

5: 2.78 M

6: 0.5 M

Part 3: The volumes are calculated:

Volumes = n / M = 55x10 ^ -6 mol / 1 M = 5.5x10 ^ -5 L

8: 2 L

9: 2x10 ^ -5 L

If you liked the answer, please rate it positively, you would help me a lot, thanks.